one of two numbers exceeds the other by 9.four times the smaller added to five times the larger gives 108. find the numbers ?

Answer by sachi(548)

let the smaller no=x
so the larger no =x+9
so 4x+5(x+9)=108
or 4x+5x+45=108
or 9x=108-45=63
or x=63/9=7
so the nos are 7 & 16
ans